### v1

# D&M: Enriching E-commerce Videos with Sound Effects by Key Moment Detection and SFX Matching

Abstract
Videos showcasing specific products are increasingly important for E-commerce. Key moments naturally exist as the first appearance of a specific product, presentation of its distinctive features, the presence of a buying link, etc. Adding proper sound effects (SFX) to these key moments, or video decoration with SFX (VDSFX), is crucial for enhancing the user engaging experience. Previous studies about adding SFX to videos perform video to SFX matching at a holistic level, lacking the ability of adding SFX to a specific moment. Meanwhile, previous studies on video highlight detection or video moment retrieval consider only moment localization, leaving moment to SFX matching untouched. By contrast, we propose in this paper D&M, a unified method that accomplishes key moment detection and moment to SFX matching simultaneously. Moreover, for the new VDSFX task we build a large-scale dataset SFX-Moment from an E-commerce platform. For a fair comparison, we build competitive baselines by extending a number of current video moment detection methods to the new task. Extensive experiments on SFX-Moment show the superior performance of the proposed method over the baselines. Code and data will be released.
1 Introduction
The eye sees better when the sound is great111by Steven Spielberg. In this paper we aim to automatically add sound effects (SFX) to key moments, which is also auto-detected, in E-commerce videos to enhance their expressiveness, see Fig. 1.
Videos showcasing specific products are increasingly important for E-commerce. Key moments naturally exist as the first appearance of a specific product, presentation of its major features, the presence of a buying link, etc. Adding proper SFX to such key moments is crucial for enhancing the engaging experience that helps turn a video viewer to a product consumer. Previous studies about adding SFX to videos either perform video to SFX matching [1, 17, 26] or video to SFX generation [6, 11, 4], both at a holistic level. They thus lack the ability of adding SFX to a specific moment. Meanwhile, previous studies on video highlight detection [14, 16] or video moment retrieval [22, 20] consider only moment localization, leaving moment to SFX matching untouched.
[…]
We go beyond video-to-SFX matching by proposing a new task termed Video Decoration with SFX (VDSFX). In particular in the context of E-commerce field, VDSFX is to automatically add one or multiple SFX to key moments in a given e-commerce video for better engaging experience. The task is thus more practical yet more challenging. As exemplified in Fig. 1, simply combining a state-of-the-art video moment detection model [20] and moment-to-SFX matching is insufficient. The questions of where in the video to add SFX and which SFX to add have to be jointly answered.
Naturally, no public data exists for the new task. To fill out the gap, we contribute SFX-Moment, a dataset for VDSFX. With over 16k highly selected videos, nearly 40k key moments and 356 distinct SFX in total, SFX-Moment is the first dataset for supporting research on VDSFX, to the best of our knowledge.
Inspired by the great success of DETR-based methods for video highlight detection  [14, 22], we propose D&M, a DETR-based method that performs key moment detection and moment-to-SFX matching simultaneously, see Fig. 2. Despite its relatively simple network architecture, training D&M is nontrivial due to the misalignment between the video and SFX representations and the inclusion of false negative examples by commonly used hard negative sampling. For more effective training, we propose moment-SFX matching (MSM) based pre-training for better initialization and tag-aware negative sampling (TaNS) to balance the hard and false negatives. In sum, our major contributions are as follows:
Our key contribution can be summarized as follows:
  • We introduce a new task, Video Decoration with SFX (VDSFX), aiming at detecting key moments in a given video and adding SFX for corresponding key moments. A task-specific dataset SFX-Moment is developed and will be released for prompting related research.
  • We propose a DETR-based method D&M, which performs key moment detection and SFX matching simultaneously. Two task-specific training strategies are developed to ensure successful network training.
  • Extensive experiments on SFX-Moment show the viability of the proposed method and its superior performance over multiple competitive baselines.
2 Related Work
As we have discussed, tackling the VDSFX task requires answering the combined question of where in the video to add SFX and which SFX to add. While a complete answer remains to be developed, partial answers exist as video highlight detection / video temporal grounding to the first part of the question and video-to-SFX retrieval to the second part. We thus briefly review progresses in the two topics.
2.1 Video Highlight Detection
Video Highlight Detection (VHD) concerns about finding out interesting or significant moments within a given video. Previous methods [25, 31, 10, 28] are ranking-based, where models attached each video clip with a salient score and retrieves clips with high scores as highlights. In addition to only utilizing the visual modality, extra data modalities are employed to the task. [2, 16] include audio modality to enhance highlight prediction, while [19] used texts to guide the selection of thumbnails.
Inspired by DETR [3], [14] proposes a DETR-based methods Moment-DETR to handle video highlight detection and video moment retrieval (VMR) given a langurage query in one architecture, since these two tasks share common characteristics of video content understanding and clip localization. After that, more DETR-based practices [29, 21, 22, 12] are conducted to improve Moment-DETR and DETR gradually becomes a paradigm in this field, while Encoder-only methods [18, 20] make competitive achievements. VDSFX shares the same moment localization target as VHD/MR. Since DETR-based methods directly localize moments, which is more efficient than Encoder-only methods where there are alternative predicted moments and complex post-processing is indispensable. Consider of that, our D&M employs a DETR architecture. Though the VHD/MR has the ability of moment localization, they have no module to perform moment-to-SFX matching, which means they can not satisfy the VDSFX. Our D&M goes further by adding moment-to-SFX matching module and conducting key moment detection and moment-to-SFX matching simultaneously.
Figure 2: Conceptual diagram of our proposed D&M method for VDSFX. The video as an example consists of $n=30$ frames with $m=9$ subtitles. “SFX0” is a special token indicating “no SFX”. The ASR module and the visual / textual / audio backbones, i.e. ViT / RoBERTa / AST, are all frozen. Non-trainable blocks are shown in gray. Best viewed on screen.
2.2 Video-to-SFX Retrieval
Video-to-SFX retrieval, a crucial component of video editing, aims to match SFX with a given video to create an engaging multimedia experience. Existing approaches can be categorized into One-Tower and Two-Tower methods. [1] introduced $L^{3}$ Net, a One-Tower model that utilizes two CNN-based feature extractors for video frames and SFX, followed by a fusion layer to determine their correspondence. Building on $L^{3}$ Net, [5] further explored various audio embeddings.
Two-Tower methods, on the other hand, focus on modality-specific encoders, particularly for audio. [26] simply projected an audio signal into the CLIP image embedding space and compared multiple audio encoders, including PANNS [13], AudioCLIP [9], and Wav2CLIP [27]. [17] assigned each sound effect a text label and utilized the CLIP text encoder for representation learning, while [7] introduced ImageBind, which uses videos or images to bind other modalities, with SFX encoded via AST [8].
However, SFX in existing Video-to-SFX Matching methods are typically encoded using a single-modal approach, which we believe is insufficient. Our D&M addresses this limitation by conducting SFX representation learning in a multi-modal manner. Moreover, existing methods perform video-to-SFX matching at a holistic level, lacking the capability to add SFX to specific moments. Our D&M model overcomes this challenge by introducing an additional key moment detection branch, transforming video-to-SFX matching into moment-to-SFX matching, which effectively meets the requirements of VDSFX.
3 SFX-Moment: A Dataset for VDSFX
To the best of our knowledge, no public dataset exists for VDSFX. To fill out the gap, we construct SFX-Moment, a set of 16,942 high-quality videos sampled from an E-commerce platform. Each video has at least one key moment. Each key moment has a meme sound effect manually added by the video creator, which we treat as ground truth. See Table 1 for an overview and more details in the supplement. Next, we detail the data construction process.
Data split Videos Total duration ASR subtitles Key moments Unique SFX
Total 16,942 215.8 hrs 398,603 39,670 356
training 13,560 173.6 hrs 320,508 31,836 356
validation 1,705 21.6 hrs 39,929 4,021 291
testing 1,677 20.6 hrs 38,166 3,813 283
Table 1: Basic statistics of our SFX-Moment dataset. Note that a specific sound effect from a given SFX set can be applied to multiple videos. We consider a closed set of SFX.
Data gathering. We initially collected approximately 100k E-commerce videos with manually added SFX from an E-commerce platform with permission. The mp3 audio files and metadata of the SFX including IDs and start-/end- times in the videos are also gathered, resulting in an initial set of 4k SFX.
Figure 3: Visualization of data statistics of SFX-Moment.
Data cleaning. We remove videos that fail to meet any of the following ad-hoc rules:
$\bullet$ 200+ clicks;
$\bullet$ 100+ follows;
$\bullet$ Duration between 3 seconds to 2 minutes;
$\bullet$ With non-rare SFX (we consider a sound effect rare if it is used fewer than 10 times).
Consequently, we preserve 16,942 videos, each with one or more SFX from a set of 356 unique SFX.
Data post-processing. Per video, we use Whisper large-v3 [23], a state-of-the-art model for automatic speech recognition (ASR), to transcribe the video narration into a sequence of timestamped sentences. Inspired by the successful use of music tags in video-to-music retrieval [24, 15], we manually categorize each sound effect to one of the following six tags: prompt, transition, humor, action, surprise and others, see Table 2. Furthermore, per sound effect, a short description is manually written to reflect its characteristics and potential use. Given the relatively small size of the SFX set, such manual annotation is affordable. As a result, each sound effect, associated with an audio clip, a short description and a categorical tag, is multi-modal. SFX-Moment will be released for reproducible research.
Tag SFX Tagging criterion
prompt 77 Attract attention when new things appear.
transition 49 Transfer focus from the current scene to the next one.
humor 53 Enhance the humor when the scene is humorous .
action 71 Dub for everyday activities.
surprise 34 Express surprise or shock.
others 72 Other SFX.
Table 2: Tags for SFX categorization.
4 Proposed D&M Method
A desirable method for VDSFX shall find temporally where in a given video to add a sound effect and what sound effect to add. More formally, we use $v$ to denote the given video, with its visual content represented by $n$ evenly sampled frames $\{f_{i}\}_{i=1}^{n}$. An off-the-shelf ASR tool shall transcribe the associated narration into a sequence of $m$ timestamped sentences $\{(b_{j},e_{j},t_{j})\}_{j=1}^{m}$, where $b_{j}$ and $e_{j}$ indicate the beginning and end timestamps, measured by frame indices, of sentence $t_{j}$. Suppose we have access to a set of $l$ SFX $\{(a_{k},d_{k},y_{k})\}_{k=1}^{l}$, with $a_{k}$ indicating the audio clip of the $k$-th sound effect, $d_{k}$ as a short description and $y_{k}$ as a categorical tag. We get around the two challenges, i.e. key moment detection and moment-SFX matching, in a joint framework, dubbed D&M. We describe the proposed network in Sec. 4.1, followed by a task-specific training strategy in Sec. 4.2.
4.1 Network
The network structure of D&M is illustrated in Fig. 2. We largely follow Moment-DETR [14], using a multi-Transformer encoder for multi-modal feature extraction and a DETR-decoder for multi-head prediction. However, as Moment-DETR is targeted at fore-/back-ground moment classification, we need to introduce multiple VDSFX-specific modules as follows.
Multi-modal video embedding. Given the video and its subtitles as a multi-modal input, we first utilize a pre-trained CLIP model [^2] Chinese CLIP [30], as our data is in Chinese. to extract $n$ frame-level and $m$ sentence-level embeddings, respectively. As a rule of thumb, each frame / sentence is represented by its CLS token. To simplify our notation, we re-use $f$ and $t$ to denote the two sorts of embeddings. Each then goes through an FC layer for domain-adapted learning. The re-learned tokens, with their positional embeddings $pe_{0}$ / $pe_{1}$ added, are concatenated to form an array of $n+m$ mixed-modal tokens denoted as $\{\hat{f}_{i}\}_{i=1}^{n}$, $\{\hat{t}_{j}\}_{j=1}^{m}$. These tokens are fed to a multi-Transformer encoder to produce $n+m$ multi-modal tokens $\{z_{i}\}_{i=1}^{m+n}$. The above process can be expressed more formally as
$\left\{\begin{array}[]{lll}\{f_{i}\}_{i=1}^{n}&\leftarrow&\text{ViT-B/32}(\{f_{i}\}_{i=1}^{n}),\\ \{t_{j}\}_{j=1}^{m}&\leftarrow&\text{RoBERTa}(\{t_{j}\}_{j=1}^{m}),\\ \{\hat{f}_{i}\}_{i=1}^{n}&\leftarrow&\text{FC}_{512\times 512}(\{f_{i}\}_{i=1}^{n})+pe_{0},\\ \{\hat{t}_{j}\}_{j=1}^{m}&\leftarrow&\text{FC}_{512\times 512}(\{t_{j}\}_{j=1}^{m})+pe_{1},\\ \{z_{i}\}_{i=1}^{n+m}&\leftarrow&\text{Transformers}(\{\hat{f}_{i}\}_{i=1}^{n},\{\hat{t}_{j}\}_{j=1}^{m}).\end{array}\right.$ (1)
Multi-modal SFX embedding. Given a sound-effect triplet $(a,d,y)$, we first employ AST of ImageBind [7], the previously used CLIP and an embedding layer to extract the audio, sentence and tag features, indicated by $a$, $d$ and $e(y)$, respectively. A multi-modal embedding of the sound effect, denoted as $s$, is obtained by a simple network as follows
$\left\{\begin{array}[]{lll}a&\leftarrow&\text{AST}(a),\\ d&\leftarrow&\text{RoBERTa}(d),\\ e(y)&\leftarrow&\text{Embedding}_{l\times 512}(y),\\ s&\leftarrow&\text{FC}_{\cdot\times 512}(\text{FC}_{\cdot\times 512}([a,d])+e(y)).\end{array}\right.$ (2)
We thus obtain the SFX embeddings as $\{s_{k}\}_{k=1}^{l}$. Note that in addition to the normal SFX, we specify a special SFX0, with its trainable embedding $s_{0}$, to handle moments that cannot be matched with any of the normal SFX.
Key moment detection and embedding. With the video and SFX embeddings available, we are now ready to invoke a DETR-decoder for key moment detection and embedding. In particular, the decoder uses an array of $p$ learnable queries $\{q_{j}\}_{j=1}^{p}$ to progressively conduct cross-attention based interaction with the video embeddings, resulting in $p$ output embeddings $\{\hat{q}_{j}\}_{j=1}^{p}$. The output is fed in parallel into two distinct MLPs, one for moment localization and the other for moment embedding. More formally, we have
$\left\{\begin{array}[]{lll}\{\hat{q}_{j}\}_{j=1}^{p}&\leftarrow&\text{DETR-Decoder}(\{q_{j}\}_{j=1}^{p},\{z_{i}\}_{i=1}^{n+m}),\\ \{pos_{j}\}_{j=1}^{p}&\leftarrow&\text{MLP}_{loc}(\{\hat{q}_{j}\}_{j=1}^{p}),\\ \{e_{j}\}_{j=1}^{p}&\leftarrow&\text{MLP}_{emb}(\{\hat{q}_{j}\}_{j=1}^{p}),\end{array}\right.$ (3)
where $pos_{j}$ indicates (normalized) start-/end-positions of the moment predicted w.r.t. the $j$-th query and $e_{j}$ is a 512-d embedding of the predicted moment.
Moment-SFX matching. Given the $l+1$ SFX embeddings and $p$ moment embeddings, a pairwise cosine similarity calcluation yields an $(l+1)\times p$ similarity matrix $M$, see Fig. 2. Column-wise max pooling over $M$ allows us to obtain the best matched sound effect per moment. Non-Maximum Suppression (NMS), with an IoU threshold of 0.3, is performed to discard lower-confidence moments. Moments matched with SFX0 are also removed. The remained moment-SFX pairs are the final output.
4.2 Task-Specific Training
Although the network structure of D&M is relatively simple, its training is nontrivial. Our early experiments show that training D&M from scratch (with the pre-trained visual / textual / audio backbones frozen) results in a suboptimal model. Our conjecture is that the SFX and the video embeddings lack semantic alignment, resulting in poor matching during training, which in turn affects the DETR-decoder and the multi-Transformer encoder. To resolve the issue, we propose moment-SFX matching (MSM) based pre-training for better initialization of the SFX encoder and the video encoder. Moreover, in order to strike a proper balance between hard and false negative samples, we propose tag-aware negative sampling (TaNS) for more efficient training.
4.2.1 Pre-training for Better Initialization
[…]
4.2.2 Tag-aware Negative Sampling
For constructing a mini batch for a given moment and sound effect pair $(u,s)$, hard negative sampling is to select from the SFX set the top similar SFX other than the ground truth. We empirically find that such a sampling strategy tends to incorrectly include either false negatives or negatives too hard to learn from. To address the issue, we propose tag-aware negative sampling. The idea is simple as follows. For a candidate negative $s^{-}$, if its tag $y(s^{-})$ is different from $y(s)$, the negative is reliable so its probability being sampled shall be proportional to its similarity to $u$, denoted as $sim(u,s^{-})$. However, if $s^{-}$ shares the same tag as $s$, the risk of being false negative is relatively high. Therefore, its sampling probability will be inversely proportional to its relative distance to the moment expressed as $|sim(u,s)-sim(u,s^{-})|$. If $sim(u,s^{-})$ is much larger than $sim(u,s)$, it is likely to be a false negative. By contrast, if $sim(u,s^{-})$ is much smaller, its value is much limited. Either case, the probability shall be reduced. More formally, we have
$\text{prob}(s^{-})\propto\left\{\begin{array}[]{cl}sim(u,s^{-}),&\text{if}~{}y(s^{-})\neq y(s),\\ -|sim(u,s)-sim(u,s^{-})|,&\text{if}~{}y(s^{-})=y(s).\end{array}\right.$ (4)
4.2.3 Loss for end-to-end training
For key moment localization, we follow Moment-DETR, computing the following two localization losses, i.e. the L1 loss and the GIoU loss, to jointly measure the misalignment between the predicted and the ground-truth moments. As for moment-SFX matching, we again compute the InfoNCE loss. The three losses are equally combined.
5 Experiments
5.1 Experimental Setup
5.1.1 Dataset
We split randomly the 16,942 videos of SFX-Moment into three disjoint subsets: 13,560 (80%) for training, 1,705 ($\sim$10%) for validation and 1,677 ($\sim$10%) for testing, see Table 1.
5.1.2 Evaluation Criteria
The evaluation of VDSFX should consider both key moment detection and corresponding SFX matching simultaneously. We employ mean Average Precision (mAP) with a predefined IoU threshold (0.5 and 0.75 in this work and by default 0.5 unless otherwise stated).In particular, we compute ${\text{AP}}_{\text{SFX}}$, which measures the ranking of predicted key moments w.r.t. a given sound effect. We obtain ${\text{mAP}}_{\text{SFX}}$ by averaging over all the test SFX. In a similar manner, we compute ${\text{AP}}_{\text{vid}}$, which measures the ranking of the matched and localized SFX w.r.t. a given test video, and ${\text{mAP}}_{\text{vid}}$ by averaging over all the test videos.
Moreover, we separately evaluate the capability of key moment detection. In addition to ${\text{AP}}_{\text{vid}}$ measuring the ranking of predicted key moments w.r.t. a given test video, we evaluate ${\text{AP}}_{\text{key}}$ which measures the ranking of predicted key moments w.r.t. a given moment class. In particular, we random choose 1k key moments in the test set and manually categorize each moment to one of the following four classes that occur frequently: introduction, transition, humour and selling, see Table 3. We obtain ${\text{mAP}}_{\text{key}}$ by averaging over the four key moment classes.
Moment class Moments Tagging criterion
introduction 532 Introduce products or their functions.
transition 134 The current scene transfers to the next one.
humour 161 The scene or the narration is humorous .
selling 173 Appeal for buying.
Table 3: Classes for key moment categorization.
5.1.3 Implementation Details
All experiments are conducted within the PyTorch framework on eight NVIDIA TESLA V100 32GB GPUs. The size of the video / moment / SFX embeddings is 512. The max length of frames and subtitles is empirically set to 20 and 40, respectively. The number of the learnable queries in DETR-Decoder is 10. We use AdamW with an initial learning rate of 3e-4 with a weight decay of 1e-4. The learning rate is decayed in a cosine schedule strategy. The MSM sub-network for pre-training is trained for 100 epochs at maximum with a mini-batch size of 512. D&M is trained for 300 epochs with a mini-batch size of 256. D&M will be open source.
5.2 Comparison with Baselines
5.2.1 Baselines
As aforementioned, no baseline method exists for VDSFX. A relatively straightforward solution is to first produce a number of candidate moments by a sliding window strategy, and then use the trained moment-to-SFX matching(MSM) sub-network to find the best matches moment-SFX pairs. We term this naive baseline Sliding-Win.
An obvious drawback of Sliding-Win is its extremely high computation cost. For more efficient key moment detection, we employ VHD methods to detect key moments before MSM. The methods should be open-source and support video frames and subtitles as input, so we choose the following representative methods that meets the requirements:
  • Two DETR-based methods: Moment-DETR [14] and QD-DETR [22], which can directly localize key moments.
  • Two encoder-only methods: UniVTG [18] and $\text{R}^{2}$-Tuning [20] that predict boundary offsets of each frame and get final key moments through NMS.
[…]
Method ${\text{mAP}}_{\text{SFX}}$ ${\text{mAP}}_{\text{vid}}$ Mean
@0.5 @0.75 @0.5 @0.75 @0.5 @0.75
Sliding-Win 9.3 3.4 31.5 14.6 20.4 9.0
[…]
D&M 12.8 6.8 34.3 21.1 23.6 14.0
Table 4: Performance of different methods for VDSFX.
Method ${\text{mAP}}_{\text{vid}}$ ${\text{mAP}}_{\text{key}}$ Mean
@0.5 @0.75 @0.5 @0.75 @0.5 @0.75
Sliding-Win 43.7 17.9 23.5 4.4 33.6 11.2
Moment-DETR [14] 31.6 12.0 16.4 3.1 24.0 7.6
QD-DETR [22] 24.1 7.0 7.2 0.6 15.7 3.8
UniVTG [18] 55.8 26.1 37.2 10.8 46.5 18.5
$\text{R}^{2}$-Tuning [20] 55.7 26.5 38.2 11.9 47.0 19.2
D&M 59.9 28.9 40.0 14.6 50.0 21.8
Table 5: Performance of different methods for Key Moment Detection.
5.2.2 Results
The performance of the baselines and the proposed D&M method for VDSFX is reported in Table 4. Among the baselines, the two DETR-based methods perform clearly worse than their Encoder-only counterparts. Sliding-Win, while simple, shows relatively descent performance when compared to the other baselines. D&M outperforms all the baselines in terms of all metrics.
The performance of the methods on the Key Moment Detection subtask is summarized in Table 5. D&M again performs consistently better than the baselines under all metrics. The effectiveness of D&M is thus verified.
5.3 Understanding D&M
Recall our D&M achieves superior performance through proposed moment-SFX matching (MSM) based pre-training and Tag-aware negative sampling (TaNS). We analyse the effectiveness and rationality of these two schemes with several ablation studies. We also analyse the necessity of the individual modalities in the video encoder and the SFX encoder. We report the ${\text{mAP}}_{\text{SFX}}$ ${\text{mAP}}_{\text{vid}}$, and their mean.
5.3.1 The Need of MSM based Pre-training
As Table 6 shows (Setup#0 vs. #Setup#1), no pre-training results in an absolute loss of 6.0 points in the overall performance.
5.3.2 The Need of TaNS
Substituting unfiorm sampling for TaNS, indicated by Setup#6, causes an absolute loss of 2.3 points. Using hard negative sampling (HardNS, Setup#7) is better than uniform sampling, but remains worse than TagNS. Moreover, we also try one-side TaNS (Setup#8), allowing only negatives that satisfy $y(s^{-})\neq y(s)$. Its marginally better performance than HardNS (22.3 vs. 22.1) suggests that excluding negatives that having the same tag as the positive SFX may help reduce false negatives. Still, TaNS is the best option.
5.3.3 MSM based Pre-training vs. TaNS
While both MSM based pre-training and TaNS are found to be necessary, the former contributes more to the good performance of D&M.
5.3.4 Alternatives to MSM for Pre-training?
We try two alternatives:
$\bullet$ Salience Score Prediction (SSP). SSP is previously used by [14] as an auxiliary task for key moment judgement. In the current context, training SSP allows us to initialize the video encoder.
$\bullet$ SFX Tag Prediction (SFX-TP). The moment embedding $u$ is fed to an MLP to predict the SFX tag of the corresponding sound effect. Such a task let the video encoder be SFX-tag aware to some extent. The clearly lower performance of SSP (Setup#2) and SFX-TP (Setup#3) suggests that initializing the video encoder alone is insufficient.
# Setup ${\text{mAP}}_{\text{SFX}}$ ${\text{mAP}}_{\text{vid}}$ Mean
0 Full 12.8 34.3 23.6
Pre-training:
1 w/o pre-training 7.6 27.6 17.6(-6.0$\downarrow$)
2 MSM $\rightarrow$ SSP 7.8 27.4 17.6(-6.0$\downarrow$)
3 MSM $\rightarrow$ SFX-TP 8.1 27.9 18.0(-5.6$\downarrow$)
When to use MSM?
4 pre-train $\rightarrow$ train 9.8 30.9 20.4(-3.2$\downarrow$)
5 pre-train $\rightarrow$ pre-train, train 12.6 33.5 23.1(-0.5$\downarrow$)
Negative Sampling:
6 TaNS $\rightarrow$ Uniform Sampling 10.2 32.4 21.3(-2.3$\downarrow$)
7 TaNS $\rightarrow$ HardNS 11.4 32.7 22.1(-1.5$\downarrow$)
8 TaNS $\rightarrow$ one-side TaNS 11.8 32.8 22.3(-1.3$\downarrow$)
Input modalities for video encoder:
9 w/o Subtitles 6.6 24.9 19.7(-3.9$\downarrow$)
10 w/o Frames 9.1 30.3 15.8(-7.8$\downarrow$)
Input modalities for SFX encoder :
11 w/o audio 8.3 26.9 17.6(-6.0$\downarrow$)
12 w/o description 10.3 30.1 20.2(-3.4$\downarrow$)
13 w/o tag 10.6 30.7 20.7(-2.9$\downarrow$)
Table 6: Ablation studies. The mAPs are computed given IoU threshold of 0.5.
Figure 4: Some qualitative results. Best viewed digitally.
5.3.5 When to use MSM?
We try two options:
$\bullet$ Option-1: Using MSM directly as an auxiliary task for end-to-end training of D&M (Setup#4)
$\bullet$ Option-2: First using MSM for pre-training and re-using it as an auxiliary task for the subsequent end-to-end training (Setup#5).
The better performance of Setup#4 against Setup#1 (20.4 vs. 17.6) shows that using MSM as an auxiliary task is beneficial when training D&M from scratch. Nevertheless, its lower performance than Setup#0 (20.4 vs 23.6) shows that MSM is best used for pre-training. The double use of MSM as Option-2 is unnecessary.
5.3.6 The influence of input modalities for video
The clearly lower performance of Setup#9 and Setup#10 shows that both frames and subtitles matter, yet the former is more important.
5.3.7 The influence of input modalities for SFX
Similar to the previous experiment, we observe that all the three modalities, i.e. audio, description and tag, are necessary for effective SFX embeddings. Meanwhile, the audio modality is the most important, followed by the description modality and the tag modality. Such a rank is consistent with the (hidden) information each modality carries about SFX.
Also, the ablation studies about the input modalities reveal the visual-audio based nature of the VDSFX task.
5.3.8 Qualitative results
Some qualitative results are shown in Fig. 4. For key moment detection, our D&M’s predictions are largely consistent with ground truth. Yet, the question of which SFX to choose remains challenging, leaving much room for future improvement.
6 Summary and Conclusions
We introduce in this paper the new task of video decoration with sound effects (VDSFX) and accordingly develop a task-specific dataset SFX-Moment. To tackle the new task, we propose D&M, an end-to-end network that performs key moment detection and moment-to-SFX matching simultaneously. Extensive experiments on the SFX-Moment dataset allow us to draw conclusions as follows. Given the challenging nature of the task, moment-SFX matching (MSM) based pre-training and tag-aware negative sampling (TaNS) are important for effective training of the D&M network. The proposed method compares favorably against multiple competitive baselines. With the developed dataset and method, our work opens up new possibilities for intelligent E-commerce video analysis and editing.